OpenStudy (anonymous):

1.Find the volume of a pyramid which has a square base of 14 inches and a height of 24 inches.

OpenStudy (anonymous):

iknoee its v= 1/3 lwh .. ?

OpenStudy (apoorvk):

Yes! actually the volume of any pyramidal form is -> 1/3 * area of its base * height

OpenStudy (anonymous):

The base is already lw, so just plug it in. V = 1/3Bh V = 1/3(14)(24) V = 112

OpenStudy (apoorvk):

so what would be the area of the square given that its edge is 14 inches long?